# Irreducible symplectic varieties with a large second Betti number

### v3

@@ 1.2. A new deformation type of irreducible symplectic varieties
As we mentioned, to verify assumptions in Theorem 1.1, we need to show that $U$ admits a compactification and the Lagrangian fibration structure can be extended over all codimension one points. Here we use the geometric invariant theory (GIT) to construct a natural concrete projective compactification $M\cong\mathbb{P}(1^{15},2^{6},3)$ of $U$ (cf. Theorem 4.8). More precisely, $M$ arises as the coarse moduli space of the Deligne–Mumford stack $\mathcal{M}$ parametrizing GIT stable pairs $(Y,H)$ with respect to the line bundle $\mathcal{O}(m,(\frac{1}{2}-\varepsilon)m)$ ($0<\varepsilon\ll 1$) on $\mathbb{P}({\rm Sym}^{3}(V^{*}))\times\mathbb{P}(V^{*})$, where $Y$ is a cubic fivefold, $H$ is a hyperplane in $\mathbb{P}^{6}$, and $Y\cap H\cong X$. To precisely characterize the moduli space $M$, we use the viewpoint of variation of GIT, which has been considered in [22] in our setting. See Sections 4.1 and 4.2 for more details of the construction and calculations. Then the extension condition is essentially implied by the results in [31], which says the Lagrangian fibration can be extended over points parametrizing cubic fivefolds $Y$ with at most one nodal singularity (we follow [40] to reprove a slightly stronger result).
@@
    The fact that the base of the Lagrangian fibration is not isomorphic to $\mathbb{P}^{21}$ suggests that $\overline{\mathcal{J}}$ is not smooth, as we expect the base of any Lagrangian fibration of an irreducible symplectic manifold is smooth and isomorphic to a projective space (this is only known in dimension $\leq 4$, see e.g. [48, 29]). However, we cannot confirm this for now.
@@ Acknowledgements
ZL would like to thank his supervisor Yongbin Ruan for encouragement and support. CX would like to thank Daniel Huybrechts, Giulia Saccà, and Richard Thomas who are co-organizers of the Oberwolfach workshop “Algebraic Geometry: Wall Crossing and Moduli Spaces, Varieties and Derived Categories”, where his interest in this topic was raised. The collaboration was started at the National Algebraic Geometry Conference organized by Shandong University, to which we would like to thank the organizers. We would like to thank Emanuele Macrì, Kieran O’Grady, and Giulia Saccà for discussions and pointing out inaccuracies in our early draft, and Giulia Saccà for informing us about her work [50]. We would like to thank Luca Tasin for pointing out a mistake in the previous version. We would also like to thank Lie Fu, Hanfei Guo, Xiaolong Liu, Songtao Ma, Mirko Mauri, Alexander Perry, Yongbin Ruan, Junliang Shen, Rui Xiong, Qizheng Yin and Xiaolei Zhao for many useful discussions on related topics.
@@ 3.1. Symplectic compactifications
Now, we prove $K_{X}\sim_{\mathbb{Q}}0$. As $\widetilde{X}\dashrightarrow X$ is the output of MMP, $K_{X}\sim_{\mathbb{Q}}D:=\rho_{*}G$ is effective. Note that if $D$ is non-zero, then it is a degenerate divisor in the sense of [38, Definition 2.9]. Indeed, if $\mathrm{codim}_{B}(\pi(D))\geq 2$, then $D$ is $\pi$-exceptional in the sense of [38, Definition 2.9]. If $\mathrm{codim}_{B}(\pi(D))=1$, as $\mathrm{codim}_{B}(B\setminus U)\geq 2$, we know that $\pi(D)$ meets $U$. Let $E$ be an irreducible component of ${\rm Supp}(D)$ whose image meets $U$. Since $X_{0}\to U$ is surjective, there exists a prime divisor $\Gamma$ of $X_{0}$ which satisfies that $\overline{\pi(\Gamma)}=\pi(E)$. Since $\Gamma\nsubseteq D$, $\Gamma\neq E$. Hence, $E$ is of insufficient fiber type in the sense of [38, Definition 2.9]. Therefore, if $D$ is non-zero, [38, Lemma 2.10] implies the diminished base locus $\mathbf{B}_{-}(K_{X}/B)\neq 0$, which contradicts the nefness of $K_{X}$ over $B$. This proves $K_{X}\sim_{\mathbb{Q}}0$.
@@ 3.2. Irreducible symplectic compactifications
Let $A$ be an abelian variety and $\pi\colon A\to B$ be a Lagrangian fibration onto a normal projective variety $B$. Then each fiber of $\pi$ is isomorphic to an abelian subvariety $A^{\prime}\subset A$.
@@
Let $T\subset U$ be the locus such that $A_{t}:=\pi^{-1}(t)$ is a simple Lagrangian torus for each $t\in T$. By assumption (3), $T$ contains a non-empty complement of a countable union of closed subvarieties of $B$, hence $T$ is dense in $B$. As
@@
such that $\pi(x)\in T$. By the density of $T\cap\operatorname{im}(\pi_{i})$ in $\operatorname{im}(\pi_{i})$, we can vary $z_{i}$ to assume that $\pi_{i}^{-1}(\pi(x))$ is a very general fiber of $\pi_{i}$ by keeping $\pi(x)\in T$. Hence Lemma 2.8 and (3) imply that $\pi_{i}^{-1}(\pi(x))$ contains an abelian variety $A^{\prime}$ of dimension $\frac{1}{2}\dim(Z_{i})$. Since $\pi_{i}^{-1}(\pi(x))\subset\pi^{-1}(\pi(x))$ and $x\in T$, we know that $A^{\prime}$ is contained in $\pi^{-1}(\pi(x))$, which is a simple abelian variety by assumption. Therefore, the only possibility is $A^{\prime}=\pi^{-1}(\pi(x))$, which implies $\dim(Z_{i})=\dim(X)$. This shows that $m=1$ and $\dim(Z_{0})=0$ as desired, and we can conclude that $X=Z_{1}$ is an irreducible symplectic variety. ∎
@@
Let $(X,\sigma)$ be a projective symplectic variety and $\pi\colon X\to B$ be a surjective morphism onto a normal projective variety $B$. Let $U\subset B$ be a smooth open subscheme with $\pi^{-1}(U)\to U$ smooth, and we assume each connected component of $\pi^{-1}(b)$ is a Lagrangian torus for a point $b\in U$. Then for any normal variety $Y$ with a finite quasi-étale covering $f\colon Y\to X$, each connected component of $(\pi\circ f)^{-1}(b)$ is a Lagrangian torus of the symplectic variety $(Y,f^{[*]}(\sigma))$ and is isogenous to a component of $\pi^{-1}(b)$.
@@
Let $(X,\sigma)$ be a projective symplectic variety and $\pi\colon X\to B$ be a surjective morphism onto a normal projective variety $B$. If connected components of very general fibers of $\pi$ are simple Lagrangian tori and $\pi$ is not isotrivial, then $(X,\sigma)$ is an irreducible symplectic variety.
@@
After taking Stein factorization, we can assume that $\pi$ is a Lagrangian fibration. Let $f\colon Y\to X$ be any finite quasi-étale morphism. To prove $(X,\sigma)$ is irreducible symplectic, it suffices to prove $(Y,f^{[*]}(\sigma))$ is a cohomological irreducible symplectic variety. From Lemma 2.4, it suffices to find an irreducible symplectic variety with a finite quasi-étale covering onto $Y$.
@@
Now we fix an arbitrary ample divisor $H$ on $X$, then we have a family of polarised abelian varieties $\pi^{-1}(V)\to V$, which induces a morphism $p\colon V\to\mathbf{A}$ to the corresponding moduli space $\mathbf{A}$ of polarised abelian varieties given by $b\mapsto(\pi^{-1}(b),H|_{\pi^{-1}(b)})$. Note that the set of isomorphism classes of abelian varieties $A^{\prime\prime}$ that $A^{\prime}$ has an isogeny of degree at most $N$ to $A^{\prime\prime}$ is finite, this means there exists a dense set of points $W\subset V$ such that $\pi^{-1}(b)\cong\pi^{-1}(b^{\prime})$ for any $b,b^{\prime}\in W$. Since $\pi^{-1}(b)$ is simple, up to translation, we can further assume $(\pi^{-1}(b),H|_{\pi^{-1}(b)})\cong(\pi^{-1}(b^{\prime}),\pi^{-1}(b)|_{\pi^{-1}(b^{\prime})})$ as polarised abelian varieties. Thus $W$ is contracted by $p$, which implies that $V$ is also contracted by $p$ as $W$ is dense in $V$. In other words, $\pi^{-1}(V)\to V$ is isotrivial and we get a contradiction. This completes the proof. ∎
@@ 4.3. The Jacobian fibration
There exists a holomorphic symplectic structure $\sigma_{V_{s}}$ on $\mathcal{J}(\mathcal{Y}_{V_{s}})$ such that each fiber of $\pi_{V_{s}}\colon\mathcal{J}(\mathcal{Y}_{V_{s}})\to V_{s}$ is Lagrangian. Moreover, $\sigma_{V_{n}}$ extends to a holomorphic $2$-form on any smooth partial compactification of $\mathcal{J}(\mathcal{Y}_{V_{s}})$.
@@
By [40, Lemma 1.1], as ${\rm CH}^{3}(Y)_{\rm hom}\to\mathcal{J}(Y)$ is surjective for a smooth cubic fivefold $Y$ (see [11]), there is a codimension 3 cycle $\mathcal{Z}\subset\mathcal{J}(\mathcal{Y}_{V_{s}})\times\mathcal{Y}_{V_{s}}$, which induces an isomorphism
@@
So $Z\circ\mathcal{Z}\in{\rm CH}^{3}(\mathcal{J}(\mathcal{Y}_{V_{s}})\times X)$. Then we define $\eta^{\prime}:=Z^{*}{\eta}\in\mathrm{H}^{2}(\mathcal{Y}_{V_{s}},\Omega_{\mathcal{Y}_{V_{s}}}^{4})$ and
@@
as $\mathrm{H}^{1}(\Omega_{Y_{t}}^{4})=\mathrm{H}^{2}(\Omega_{Y_{t}}^{4})=0$, $\eta^{\prime\prime}$ determines a non-zero class $\widetilde{\eta}_{t}\in\mathrm{H}^{1}(Y_{t},\Omega^{4}_{Y_{t}}(1))$. By a similar argument as in [40, Theorem 1.2], we see that $\lrcorner{\sigma_{t}}$ is given by the multiplication with the class $\widetilde{\eta}_{t}$. Then it is an isomorphism by the following lemma.

### v5

@@ 1.2. A new deformation type of irreducible symplectic varieties
As we mentioned, to verify assumptions in Theorem 1.1, we need to show that $U$ admits a compactification and the Lagrangian fibration structure can be extended over all codimension one points. Here we use the geometric invariant theory (GIT) to construct a natural concrete projective compactification $M\cong\mathbb{P}(1^{15},2^{6},3)$ of $U$ (cf. Theorem 4.8). More precisely, $M$ arises as the coarse moduli space of the Deligne–Mumford stack $\mathcal{M}$ parametrizing GIT stable pairs $(Y,H)$ with respect to the line bundle $\mathcal{O}(m,(\frac{1}{2}-\epsilon)m)$ ($0<\epsilon\ll 1$) on $\mathbb{P}({\rm Sym}^{3}(V^{*}))\times\mathbb{P}(V^{*})$, where $Y$ is a cubic fivefold, $H$ is a hyperplane in $\mathbb{P}^{6}$, and $Y\cap H\cong X$. To precisely characterize the moduli space $M$, we use the viewpoint of variation of GIT, which has been considered in [22] in our setting. See Sections 4.1 and 4.2 for more details of the construction and calculations. Then the extension condition is essentially implied by the results in [31], which says the Lagrangian fibration can be extended over points parametrizing cubic fivefolds $Y$ with at most one nodal singularity (we follow [40] to reprove a slightly stronger result).
@@
    The fact that the base of the Lagrangian fibration is not isomorphic to $\mathbb{P}^{21}$ suggests that $\overline{\mathcal{J}}$ is not smooth, as we expect the base of any Lagrangian fibration of an irreducible symplectic manifold is smooth and isomorphic to a projective space (this is only known in dimension $\leq 4$, see e.g. [48, 29]). However, we cannot confirm this for now.  [^2] After our paper was published, János Kollár pointed out to us that $\overline{\mathcal{J}}$ is indeed singular. Let $(x\in X)\to\mathbb{P}(1^{15},2^{6},3)$ be a germ of a singularity on $\mathbb{P}(1^{15},2^{6},3)$. So there is a Galois quasi-étale cover $(x^{\prime}\in X^{\prime})\to(x\in X)$ with Galois group $G=\mathbb{Z}/2\mathbb{Z}$ or $\mathbb{Z}/3\mathbb{Z}$. If $\overline{\mathcal{J}}$ is smooth, then by [29, Lemma 2.1], $\pi$ can be lifted to a $G$-equivariant fibration $\pi^{\prime}\colon\mathcal{J}^{\prime}:=\overline{\mathcal{J}}\times_{\mathbb{P}(1^{15},2^{6},3)}X^{\prime}\to X^{\prime}$ between smooth projective $G$-varieties such that the action of $G$ on $\mathcal{J}^{\prime}$ is free. Here, we use $\overline{\mathcal{J}}$ is smooth so $\mathcal{J}^{\prime}\to\overline{\mathcal{J}}$ is étale by purity of branch locus. In particular, $\pi^{-1}(x)=|G|\cdot\pi^{-1}(x)_{\rm red}$ as cycle classes. However, $\pi$ admits a rational section $Z$ with $Z\cdot\pi^{-1}(x)=1$, which is a contradiction.
@@ Acknowledgements
ZL would like to thank his supervisor Yongbin Ruan for encouragement and support. CX would like to thank Daniel Huybrechts, Giulia Saccà, and Richard Thomas who are co-organizers of the Oberwolfach workshop “Algebraic Geometry: Wall Crossing and Moduli Spaces, Varieties and Derived Categories”, where his interest in this topic was raised. The collaboration was started at the National Algebraic Geometry Conference organized by Shandong University, to which we would like to thank the organizers. We would like to thank Emanuele Macrì, Kieran O’Grady, and Giulia Saccà for discussions and pointing out inaccuracies in our early draft, and Giulia Saccà for informing us about her work [50]. We would like to thank Luca Tasin for pointing out a mistake in the previous version. We would also like to thank Lie Fu, Hanfei Guo, Xiaolong Liu, Songtao Ma, Mirko Mauri, Alexander Perry, Yongbin Ruan, Junliang Shen, Rui Xiong, Qizheng Yin and Xiaolei Zhao for many useful discussions on related topics. We thank the anonymous referee for a careful reading as well as a list of useful suggestions that improved the exposition of the paper, especially the simplification of the proof of Theorem 4.17.
@@ 3.1. Symplectic compactifications
Now, we prove $K_{X}\sim_{\mathbb{Q}}0$. As $\rho\colon\widetilde{X}\dashrightarrow X$ is the output of MMP, $K_{X}\sim_{\mathbb{Q}}D:=\rho_{*}G$ is effective. Note that if $D$ is non-zero, then it is a degenerate divisor in the sense of [38, Definition 2.9]. Indeed, if $\mathrm{codim}_{B}(\pi(D))\geq 2$, then $D$ is $\pi$-exceptional in the sense of [38, Definition 2.9]. If $\mathrm{codim}_{B}(\pi(D))=1$, as $\mathrm{codim}_{B}(B\setminus U)\geq 2$, we know that $\pi(D)$ meets $U$. Let $E$ be an irreducible component of ${\rm Supp}(D)$ whose image meets $U$. Since $X_{0}\to U$ is surjective, there exists a prime divisor $\Gamma$ of $X_{0}$ which satisfies that $\overline{\pi(\Gamma)}=\pi(E)$. Since $\Gamma\nsubseteq D$, $\Gamma\neq E$. Hence, $E$ is of insufficient fiber type in the sense of [38, Definition 2.9]. Therefore, if $D$ is non-zero, [38, Lemma 2.10] implies the diminished base locus $\mathbf{B}_{-}(K_{X}/B)\neq 0$, which contradicts the nefness of $K_{X}$ over $B$. This proves $K_{X}\sim_{\mathbb{Q}}0$.
@@ 3.2. Irreducible symplectic compactifications
Let $A$ be an abelian variety with a holomorphic symplectic form $\sigma$ and $\pi\colon A\to B$ be a Lagrangian fibration onto a normal projective variety $B$. Then each fiber of $\pi$ is isomorphic to an abelian subvariety $A^{\prime}\subset A$.
@@
Let $T\subset B$ be the locus such that $A_{t}:=\pi^{-1}(t)$ is a simple Lagrangian torus for each $t\in T$. By assumption (3), $T$ contains a non-empty complement of a countable union of closed subvarieties of $B$, hence $T$ is dense in $B$. As
@@
such that $\pi(x)\in T$. Since $T$ is a countable intersection of open dense subsets in $B$ and $T\cap{\rm im}(\pi_{i})$ is non-empty, we see that $T\cap\operatorname{im}(\pi_{i})$ is a countable intersection of open dense subsets in the variety $\operatorname{im}(\pi_{i})$ (since we work over the uncountable field $\mathbb{C}$). Therefore, we can change $z_{i}$ (but fix $z_{j}$ ($j\neq i$)) to assume that $\pi_{i}^{-1}(\pi(x))$ is a very general fiber of
$\pi_{i}\colon Z_{i}\times\prod^{m}_{j\neq i}\{z_{j}\}\to{\rm Im}(\pi_{i})\subset B$
by keeping $\pi(x)\in T$. Hence Lemma 2.8 and (3) imply that $\pi_{i}^{-1}(\pi(x))$ contains an abelian variety $A^{\prime}$ of dimension $\frac{1}{2}\dim(Z_{i})$. Since $\pi_{i}^{-1}(\pi(x))\subset\pi^{-1}(\pi(x))$ and $x\in T$, we know that $A^{\prime}$ is contained in $\pi^{-1}(\pi(x))$, which is a simple abelian variety by assumption. Therefore, the only possibility is $A^{\prime}=\pi^{-1}(\pi(x))$, which implies $\dim(Z_{i})=\dim(X)$. This shows that $m=1$ and $\dim(Z_{0})=0$ as desired, and we can conclude that $X=Z_{1}$ is an irreducible symplectic variety.
@@
Let $(X,\sigma)$ be a projective symplectic variety and $\pi\colon X\to B$ be a surjective morphism onto a normal projective variety $B$. Let $U\subset B$ be a smooth open dense subscheme with $\pi^{-1}(U)\to U$ smooth, and we assume each connected component of $\pi^{-1}(b)$ is a Lagrangian torus for a point $b\in U$. Then for any normal variety $Y$ with a finite quasi-étale covering $f\colon Y\to X$, each connected component of $(\pi\circ f)^{-1}(b)$ is a Lagrangian torus of the symplectic variety $(Y,f^{[*]}(\sigma))$ and is isogenous to a component of $\pi^{-1}(b)$.
@@
Let $(X,\sigma)$ be a projective symplectic variety and $\pi\colon X\to B$ be a surjective morphism onto a normal projective variety $B$ with general fibers Lagrangian. If connected components of very general fibers of $\pi$ are simple abelian vareities and $\pi$ is not isotrivial, then $(X,\sigma)$ is an irreducible symplectic variety.
@@
After taking Stein factorization, we can assume that $\pi$ is a Lagrangian fibration. Let $f\colon Y\to X$ be any finite quasi-étale morphism. To prove $(X,\sigma)$ is irreducible symplectic, it suffices to prove $(Y,f^{[*]}(\sigma))$ is a cohomological irreducible symplectic variety. From Lemma 2.4, it suffices to find a cohomological irreducible symplectic variety with a finite quasi-étale covering onto $Y$.
@@
Now we fix an arbitrary ample divisor $H$ on $X$, then we have a family of polarised abelian varieties $\pi^{-1}(V)\to V$, which induces a morphism $p\colon V\to\mathbf{A}$ to the corresponding moduli space $\mathbf{A}$ of polarised abelian varieties given by $b\mapsto(\pi^{-1}(b),H|_{\pi^{-1}(b)})$. Note that the set of isomorphism classes of abelian varieties $A^{\prime\prime}$ that $A^{\prime}$ has an isogeny of degree at most $N$ to $A^{\prime\prime}$ is finite, this means there exists a dense set of points $W\subset V$ such that $\pi^{-1}(b)\cong\pi^{-1}(b^{\prime})$ for any $b,b^{\prime}\in W$. Since $\pi^{-1}(b)$ is simple, up to translation, we can further assume $(\pi^{-1}(b),H|_{\pi^{-1}(b)})\cong(\pi^{-1}(b^{\prime}),H|_{\pi^{-1}(b^{\prime})})$ as polarised abelian varieties. Thus $W$ is contracted by $p$, which implies that $V$ is also contracted by $p$ as $W$ is dense in $V$. In other words, $\pi^{-1}(V)\to V$ is isotrivial and we get a contradiction. This completes the proof. ∎
@@ 4.3. The Jacobian fibration
There exists a holomorphic symplectic structure $\sigma_{V_{s}}$ on $\mathcal{J}(\mathcal{Y}_{V_{s}})$ such that each fiber of $\pi_{V_{s}}\colon\mathcal{J}(\mathcal{Y}_{V_{s}})\to V_{s}$ is Lagrangian. Moreover, $\sigma_{V_{s}}$ extends to a holomorphic $2$-form on any smooth partial compactification of $\mathcal{J}(\mathcal{Y}_{V_{s}})$.
@@
By [40, Lemma 1.1], as ${\rm CH}^{3}(Y)_{\rm hom}\to\mathcal{J}(Y)$ is surjective for a smooth cubic fivefold $Y$ (see [11]), there is a codimension 3 cycle $\mathcal{Z}\in{\rm CH}^{3}(\mathcal{J}(\mathcal{Y}_{V_{s}})\times_{V_{s}}\mathcal{Y}_{V_{s}})_{\mathbb{Q}}$, which induces an isomorphism
@@
So $Z\circ\mathcal{Z}\in{\rm CH}^{3}(\mathcal{J}(\mathcal{Y}_{V_{s}})\times X)_{\mathbb{Q}}$. Then we define $\eta^{\prime}:=Z^{*}{\eta}\in\mathrm{H}^{2}(\mathcal{Y}_{V_{s}},\Omega_{\mathcal{Y}_{V_{s}}}^{4})$ and
@@
as $\mathrm{H}^{1}(\Omega_{Y_{t}}^{4})=\mathrm{H}^{2}(\Omega_{Y_{t}}^{4})=0$, $\eta^{\prime\prime}$ determines a unique non-zero class $\widetilde{\eta}_{t}\in\mathrm{H}^{1}(Y_{t},\Omega^{4}_{Y_{t}}(1))$. As in [40, Theorem 1.2], we have the following description of $\lrcorner{\sigma_{t}}$.
Lemma 4.13.
The map $\lrcorner{\sigma_{t}}\colon T_{t,V_{s}}\cong\mathrm{H}^{1}(Y_{t},T_{Y_{t}}(-X))\to\mathrm{H}^{2}(Y_{t},\Omega^{3}_{Y_{t}})$ identifies with the multiplication with the class $\widetilde{\eta}_{t}$.
Proof.
The argument is similar to the proof of [40, Theorem 1.2]. From the exact sequence
(5) $0\to\Omega_{V_{s},t}\otimes\mathcal{O}_{Y_{t}}\to\Omega_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}\to\Omega_{Y_{t}}\to 0,$
we have an exact sequence
$0\to\Omega_{V_{s},t}\otimes\Omega^{3}_{Y_{t}}\to\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}\to\Omega^{4}_{Y_{t}}\to 0,$
where $L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}$ is a subbundle of $\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}$. Therefore, we get a commutative diagram
00{0}00{0}00{0}00{0}${{\Omega_{V_{s},t}\otimes\Omega^{3}_{Y_{t}}}}$${{\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}}}$${{\Omega^{4}_{Y_{t}}}}$00{0}00{0}${{\Omega_{V_{s},t}\otimes\Omega^{3}_{Y_{t}}(X)}}$${{(\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}})(X)}}$${{\Omega^{4}_{Y_{t}}(X)}}$00{0}00{0}${{\Omega_{V_{s},t}\otimes\Omega^{3}_{Y_{t}}(X)|_{X}}}$${{(\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{X}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{X})(X)}}$${{\Omega^{4}_{Y_{t}}(X)|_{X}}}$00{0}00{0}00{0}00{0}$\scriptstyle{f}$$\scriptstyle{g}$$\scriptstyle{k}$$\scriptstyle{h}$
with all columns and rows exact. Since the composition $X\times V_{s}\cong Z\hookrightarrow\mathcal{Y}_{V_{s}}\times X\to\mathcal{Y}_{V_{s}}$ is an embedding of a divisor, then $Z^{*}$ is given by
$Z^{*}\colon\mathrm{H}^{1}(X,\Omega^{3}_{X})\xrightarrow{p_{X}^{*}}\mathrm{H}^{1}(Z,\Omega^{3}_{Z})\to\mathrm{H}^{2}(\mathcal{Y}_{V_{s}},\Omega^{4}_{\mathcal{Y}_{V_{s}}})$
where $p_{X}^{*}\colon Z\to X$ is the projection, and the restriction map to $Y_{t}$
$\mathrm{H}^{1}(X,\Omega^{3}_{X})\xrightarrow{Z^{*}}\mathrm{H}^{2}(\mathcal{Y}_{V_{s}},\Omega^{4}_{\mathcal{Y}_{V_{s}}})\to\mathrm{H}^{2}(Y_{t},\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}})$
can be factored as
$\mathrm{H}^{1}(X,\Omega^{3}_{X})\xrightarrow{p_{X}^{*}}\mathrm{H}^{1}(Z,\Omega^{3}_{Z})\to\mathrm{H}^{1}(X,\Omega^{3}_{Z}|_{X})$
$\xrightarrow{\ell_{1}}\mathrm{H}^{1}(X,\Omega^{4}_{\mathcal{Y}_{V_{s}}}(X)|_{X})\xrightarrow{\delta_{1}}\mathrm{H}^{2}(Y_{t},\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}).$
Here, $\ell_{1}$ is the map $\mathrm{H}^{1}(X,\Omega^{3}_{Z}|_{X})\to\mathrm{H}^{1}(X,\Omega^{4}_{\mathcal{Y}_{V_{s}}}(X)|_{X})$ induced by the exact sequence
$0\to\Omega^{3}_{Z}|_{X}\to\Omega^{4}_{\mathcal{Y}_{V_{s}}}(X)|_{X}\to\Omega^{4}_{Z}(X)|_{X}\to 0$
associated with
$0\to\operatorname{\mathcal{O}}_{X}(-X)\to\Omega_{\mathcal{Y}_{V_{s}}}|_{X}\to\Omega_{Z}|_{X}\to 0$
and $\delta_{1}$ is the connecting map in the cohomology long exact sequence of
$0\to\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}\to\Omega^{4}_{\mathcal{Y}_{V_{s}}}(X)|_{Y_{t}}\to\Omega^{4}_{\mathcal{Y}_{V_{s}}}(X)|_{X}\to 0$
In particular, the restriction map
$\ell_{2}\colon\mathrm{H}^{1}(X,\Omega^{3}_{X})\to\mathrm{H}^{2}(Y_{t},\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}})$
factors through
$\mathrm{H}^{1}(X,\Omega^{3}_{X})\to\mathrm{H}^{1}(X,\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{X}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{X})\xrightarrow{\delta_{2}}\mathrm{H}^{2}(Y_{t},\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}),$
where $\delta_{2}$ is the connecting map in the cohomology long exact sequence of the middle column in the above diagram. Therefore, the class $\eta^{\prime\prime}\in\mathrm{H}^{1}(Y_{t},\Omega^{4}_{Y_{t}}(1)|_{X})$ constructed above have a lift $\eta^{\prime\prime\prime}\in\mathrm{H}^{1}(X,\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{X}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{X})$ such that
$\hat{\eta}_{t}:=\ell_{2}(\eta)=\delta_{2}(\eta^{\prime\prime\prime})\in\mathrm{H}^{2}(Y_{t},\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}/L^{2}\Omega^{4}_{\mathcal{Y}_{V_{s}}}|_{Y_{t}}).$
Since $\mathrm{H}^{4,2}(Y_{t})=\mathrm{H}^{4,1}(Y_{t})=0$, the class $\hat{\eta}_{t}$ can be lifted to a unique class
$\mathrm{int}(\cdot)Z^{*}\eta\in\mathrm{H}^{2}(Y_{t},\Omega_{V_{s},t}\otimes\Omega^{3}_{Y_{t}})=\operatorname{Hom}(T_{V_{s},t},\mathrm{H}^{2}(Y_{t},\Omega^{3}_{Y_{t}}))$
which is also regarded as a linear map. Then as in the proof of [40, Theorem 1.2], from the above construction, we have
$v\lrcorner{\sigma_{t}}=([\mathcal{Z}^{2,2}]|_{\mathcal{J}(Y_{t})\times Y_{t}})^{*}(\mathrm{int}(v)Z^{*}\eta)$
for any $v\in T_{V_{s},t}$, where
$([\mathcal{Z}^{2,2}]|_{\mathcal{J}(Y_{t})\times Y_{t}})^{*}\colon\mathrm{H}^{2}(Y_{t},\Omega^{3}_{Y_{t}})\to\mathrm{H}^{0}(\mathcal{J}(Y_{t}),\Omega_{\mathcal{J}(Y_{t})})$
is the natural isomorphism induced by $\mathcal{Z}$. Therefore, to prove this lemma, it suffices to prove that $\mathrm{int}(\cdot)Z^{*}\eta$ is the same as the multiplication with the class $\widetilde{\eta}_{t}$. This can be deduced from a diagram chasing in the cohomology long exact sequences associated with the above commutative diagram, since from the construction, we have $k(\widetilde{\eta}_{t})=\eta^{\prime\prime}=h(\eta^{\prime\prime\prime})$ and $\mathrm{int}(\cdot)Z^{*}\eta=f^{-1}(\delta_{2}(\eta^{\prime\prime\prime}))=f^{-1}(\hat{\eta})$, so $g(\mathrm{int}(\cdot)Z^{*}\eta)=\delta_{3}(\widetilde{\eta}_{t})\in\Omega_{V_{s},t}\otimes\mathrm{H}^{2}(Y_{t},\Omega^{3}_{Y_{t}}(X))$, where
$\delta_{3}\colon\mathrm{H}^{1}(Y_{t},\Omega^{4}_{Y_{t}}(X))\to\Omega_{V_{s},t}\otimes\mathrm{H}^{2}(Y_{t},\Omega_{Y_{t}}^{3}(X))$
is the connecting map associated with the second row in the above diagram. Now the result follows from $\mathrm{H}^{1}(X,\Omega^{3}_{Y_{t}}(X)|_{X})=0$. ∎
Therefore, the map $\lrcorner{\sigma_{t}}$, or equivalently, the multiplication with the class $\widetilde{\eta}_{t}$, is an isomorphism by the following lemma.